An electric train develops a power of 1.0 MW when travelling at a constant speed of 50 ms-1. The net resistive force acting on the train is (Ans: 20 kN).


My work: P=FV, F=P/V, F=2x10-6

1
Expert's answer
2022-04-25T13:33:04-0400

"P=Fv,\implies"

"F=\frac Pv=\frac{10^6}{50}=20~kN."